# Weighted Envy Freeness With Limited Subsidies

### v1

@@ 3.1 A given allocation: the weighted-envy graph
When trying to extend Theorem 2.3 to the weighted setup, a notable distinction arises: in the unweighted context, an allocation $A$ is EF-able if and only if $A$ maximizes the utilitarian welfare across all reassignments of its bundles to agents (part (2) of Theorem 2.3).
@@
Contrastingly, in the weighted setup, this assertion does not necessarily hold, as demonstrated by Example 1.1. Moreover, the converse of (3) in Theorem 2.3 can fail too.
Proposition 3.1.
There exists an allocation $A$, which is WEF-able, yet maximizes neither the utilitarian social welfare or the weighted utilitarian social welfare.
@@
It turns out that two out of three parts of this theorem still hold in the weighted setting; The proof is similar to the one in [15].
Theorem 3.2.
For an allocation $A$, the following statements are equivalent:
@@
$=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}+\frac{s_{j}}{w_{j}}$
@@
Theorem 3.2 presents an effective method for verifying whether a given allocation $A$ is WEF-able. We use it to present an alternative proof that part 2 of Theorem 2.3 does not hold in the weighted setting.
Proposition 3.3.
@@
As in Example 1.1, consider 2 bundles $B_{1},B_{2}$ and 2 agents, $N=\{i_{1},i_{2}\}$, with weights $w_{1}=1,w_{2}=10$ and valuation functions
@@
By Proposition 3.2, both permutations are not WEF-able. ∎
Deciding that an allocation $A$ is WEF-able is not enough. As in Halpern and Shah [15], we also need to find a minimal subsidy vector, $\mathbf{s}$, that ensures WEF for $(A,\mathbf{s})$. This subsidy vector can be computed in strongly polynomial time and matches the one constructed in the proof of Proposition 3.2.
Theorem 3.4.
For any allocation $A$ that is WEF-able, let $\mathbf{s}^{*}$ be a subsidy vector defined by $\mathbf{s}^{*}_{i}:=w_{i}cost_{A}(\ell_{i}(A))$, for all $i\in N$. Then
@@
    $\Rightarrow\frac{s_{i}}{w_{i}}\geq\frac{s_{i}^{*}}{w_{i}}+\frac{s_{i_{r}}}{w_{i_{r}}}\geq\frac{s_{i}^{*}}{w_{i}}$
@@
    The computation of $s^{*}$ can be executed as follows: Initially, the Floyd-Marshall algorithm ([28], [29]) is applied to the graph derived by negating all edge costs in $G_{A,w}$ (This has a linear time solution since there are no cycles with positive costs in the graph). Hence, determining the longest path cost between any two agents, accomplished in $O(nm+n^{3})$ time. Subsequently, the longest path starting at each agent is identified in $O(n^{2})$ time.
@@
    Assume, for the sake of contradiction, that $s^{*}_{i}>0$ for every agent $i\in N$, which implies that $\text{cost}_{A}(\ell_{i}(A))>0$ since $w_{i}\geq 1$. Because the number of agents is finite, there must be $k\geq 2$ agents, say $i_{1},\dots,i_{k}$, such that $\ell_{i_{1}}(A),\dots,\ell_{i_{k}}(A)$ form a cycle with positive cost in $G_{w,A}$, contradicting Theorem 3.2.
@@
Now we can find the minimum subsidy needed in the worst-case scenario for agents with different entitlements, whether the allocation is given or can be chosen. The proofs extend those of Halpern and Shah [15]. Note that Theorem 3.5 applies to any additive valuations, including identical and binary cases.
Theorem 3.5.
For every weight vector, and for every given WEF-able allocation $A$, there exists an envy-eliminating subsidy vector with total subsidy at most $\left(\frac{W}{w_{1}}-1\right)mV$, and this bound is tight in the worst case.
@@
For the lower bound, consider an instance with $m$ identical items, which all $n$ agents value at $V$. Consider the allocation $A$, which assigns all items to a single agent $j$ with the minimum entitlement. It’s evident that $A$ is WEF-able, and its optimal subsidy vector $\mathbf{s}$ satisfies $s_{j}=0$ and $s_{i}=\frac{w_{i}}{w_{j}}mV=\frac{w_{i}}{w_{1}}mV$ for $i\neq j\in N$. Therefore, we require
$\frac{W-w_{1}}{w_{1}}mV=\left(\frac{W}{w_{1}}-1\right)mV$
To establish the upper bound, the required subsidy amount is the summation of the maximum cost paths originating from different agents (Theorem 3.4). To bound the highest cost of a path starting from $i$, we make the following lemma:
Lemma 3.6.
For every WEF-able allocation $A$ and agent $i$, the highest cost of a path from $i$ in $G_{A,\mathbf{w}}$ is at most $\frac{mV}{w_{1}}$.
By Lemma 3.6, the highest cost of a path from $i$ in $G_{A,\mathbf{w}}$ is at most $\frac{mV}{w_{1}}$. So the subsidy of agent $i$ does not exceed $w_{i}\frac{mV}{w_{1}}$. By part (4) of 3.4, at least one agent has a subsidy of 0. This implies a total subsidy of at most
$\frac{W-w_{1}}{w_{1}}mV=\left(\frac{W}{w_{1}}-1\right)mV\qed$
Proof.
(Lemma 3.6) Since $A$ is WEF-able, there are no positive-weight cycles in $G_{A,w}$, so the highest cost of a path is finite.
For every two agents $i,j$,
$cost_{A}(i,j)=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq\frac{\sum_{o\in A_{j}}v_{i}(o)}{w_{1}}\leq\frac{\sum_{o\in M}v_{i}(o)}{w_{1}}$
Therefore, the cost of every path is at most $\frac{mV}{w_{1}}$. ∎
Importantly, for a fixed allocation, the weighted setting maximum worst-case subsidy depends on the proportion of $w_{1}$ in the total weight $W$, which can be much larger than the number of agents $n$, the argument that the unweighted setting depends on.
Halpern and Shah proved in [15] that, if an allocation $A$ is EF-able and EF1, then every path in $G_{A,w}$ has cost at most $(n-1)V$, and therefore a subsidy of at most $(n-1)^{2}V$ is sufficient. The following theorem generalizes this result to the weighted setting.
Theorem 3.7.
Let $A$ be both WEF-able and $WEF(x,y)$ for any $x,y\in[0,1]$ such that $x+y\leq 1$. Then there exists an envy-eliminating subsidy vector with total subsidy at most $\left(\frac{W}{w_{1}}-1\right)(n-1)V$.
@@
$\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq\frac{yv_{i}(B)}{w_{i}}+\frac{xv_{i}(B)}{w_{j}}=\frac{(1-x)v_{i}(B)}{w_{i}}+\frac{xv_{i}(B)}{w_{j}}\leq$
$\frac{(1-x)v_{i}(B)}{w_{1}}+\frac{xv_{i}(B)}{w_{1}}=\frac{(1-x)v_{i}(B)+xv_{i}(B)}{w_{1}}=\frac{v_{i}(B)}{w_{1}}\leq\frac{V}{w_{1}}$
Any path contains at most $n-1$ arcs, that is, $s_{i}=w_{i}cost_{A}(\ell_{i}(A))\leq\frac{w_{i}}{w_{1}}(n-1)V$. From part (4) of Theorem 3.4, there is at least one agent that requires no subsidy, so the required total subsidy is at most $\frac{W-w_{1}}{w_{1}}(n-1)V=\left(\frac{W}{w_{1}}-1\right)(n-1)V$.
@@ 3.2 Computing a WEF-able allocation
We start with a lower bound on the required subsidy for general additive valuations. To prove the lower bound, we need a lemma.
Lemma 3.8.
Suppose there are $n$ agents, and only one item $o$ which the agents value positively. Then an allocation is WEF-able iff $o$ is given to an agent $i$ with the highest $v_{i}(o)$.
@@
By Theorem 3.2, it is sufficient to check the cycles in the weighted envy-graph. If $o$ is given to $i$, then $i$’s envy is $-\frac{v_{i}(o)}{w_{i}}$, and the envy of every other agent $j$ in $i$ is $\frac{v_{j}(o)}{w_{i}}$. All other envies are 0. The only potential positive-weight cycles are cycles of length 2 involving agent $i$. The weight of such a cycle is positive iff $\frac{v_{j}(o)}{w_{i}}-\frac{v_{i}(o)}{w_{i}}>0$, which holds iff $v_{j}(o)>v_{i}(o)$. Therefore, there are no positive-weight cycles iff $v_{i}(o)$ is maximum. ∎
Theorem 3.9.
For every weight vector $\mathbf{w}$ and any $n\geq 2$, no algorithm can guarantee a total subsidy smaller than $(\frac{W}{w_{1}}-1)V$.
@@
By Lemma 3.8, the only WEF-able allocation is to give $o$ to agent $1$. In this case, the minimum subsidy is $s_{1}=0$ and $s_{i}=w_{i}\frac{v_{i}(o)}{w_{1}}=w_{i}V_{2nd}$. Summing all subsidies leads to $\sum_{i\geq 2}w_{i}V_{2nd}=(W-1)V_{2nd}$. As $\epsilon$ can be arbitrarily small, we get a lower bound of $(W-1)V$. ∎
Our task is to construct a WEF-able allocation $A$ with maximum subsidy of $w_{i}V$ per agent $i\in N$. In other words, the maximum cost of a path needs to be below $V$.
For the unweighted setting, Brustle et al. [7] presented an algorithm, based on iterative maximum matching between agents and items, that finds an envy-freeable allocation for which the required subsidy is at most $(n-1)V$, which is tight.
However, this algorithm might result in an allocation that is not WEF-able, as demonstrated in Example 1.1: The allocation $A_{1}=\{o_{1}\},A_{2}=\{o_{2}\}$ is derived from maximum matching between the agents and items but is not WEF-able. Intuitively, since agent 1 has greater entitlements, he should receive a more valuable bundle.
Our algorithm addresses this by finding a maximum matching between several copies of the agents, based on their entitlements, and the items.
Since our algorithm operates with copies of agents according to their entitlements, we assume the entitlements are integers. In other words, the algorithm constructs a WEF-compatible allocation for agents with integer weights.
For simplicity, if the number of items at the start of a round is less than $W$, we add dummy items (valued at 00 by all agents) such that the total number of items is at least $W$. This implies that, in each round, every agent $i$ gets exactly $w_{i}$ items (each agent can receive at most $w_{i}$ items since they have $w_{i}$ copies. Because the algorithm computes a maximum matching between the $W=\Sigma_{i\in N}w_{i}$ items and the agents’ copies, each agent must receive $w_{i}$ items).
In example 1.1, we add 9 dummy items, and then find a maximum-weight matching between $w_{1}=10$ copies of agent 1 and $w_{2}=1$ copy of agent 2 with the items. The first round of maximum weight matching gives a real item to two copies of agent 1, resulting in the allocation $A_{1}=\{o_{1},o_{2}\},A_{2}=\emptyset$, which is WEF-able.
We denote by $H[I,J]$ the complete bipartite graph on vertex sets $I$ and $J$, where edge $(i,o_{j})$ has the value $v_{i}(o_{j})$.
In round $t$, the algorithm finds a maximum-value matching $\mu^{t}$ in $H[I,O_{t}]$, where $I$ is the set of copies of the agents and $O_{t}$ consists of the items that have not yet been allocated in this round. If the $w_{i}$ copies of agent $i$, denoted $i_{1},...,i_{w_{i}}$ are matched to items $\mu_{i_{1}}^{t},...,\mu_{i_{w_{i}}}^{t}$, respectively, then the algorithm allocates these items to agent $i$. We denote the total set of items allocated to $i$ as $A_{i}^{t}=\{\mu_{i_{1}}^{t},...,\mu_{i_{w_{i}}}^{t}\}$ and say that the algorithm computed the allocation $A^{t}$ at iteration $t$.
This process continues until iteration $T$, at which all items have been allocated.
@@
Algorithm 1 Weighted Sequence Protocol For Additive Valuations
1:Instance $(N,M,v,\{w_{i}\}_{i\in N})$ with additive valuations.
2:WEF-able allocation $A=(A_{i})_{i\in N}$ with total required subsidy of at most $(W-1)V$.
@@
6:while $O_{t}\neq\emptyset$ do
7:     Compute a maximum-value matching $\mu^{t}=\{(i,\mu_{i}^{t})\}_{i\in I}$ in $H[I,O_{t}]$
8:     $A_{i}^{t}\leftarrow\{\mu_{i_{1}}^{t},...,\mu_{i_{w_{i}}}^{t}\}$, for all $i\in N$
9:     Set $A_{i}\leftarrow A_{i}\cup A_{i}^{t}$, for all $i\in N$
[…]
11:     $t\leftarrow t+1$
12:end while
13:return $A$
First, we prove that the allocation $A$ output from Algorithm 1 is WEF-able. By Theorem 3.2, it is sufficient to prove that its weighted envy-graph has no positive-cost cycles.
We prove that this holds not only for the final allocation, but also for each allocation $A^{t}$ computed in iteration $t$.
For each cycle $C$ in the envy-graph of $A^{t}$, we construct another allocation by choosing one item uniformly at random from the bundle of each agent in $C$, and transferring this item to the following agent in $C$.
We show that the expected change in value between the new allocation and the allocation returned by Algorithm 1 is at most 00; we further show that the expected change is equal to the cost of $C$. Therefore, the cost of $C$ is at most 00.
Proposition 3.10.
@@
We prove that, in every round $t$, the total cost added to any directed cycle in the weighted-envy graph is non-positive. Combined with Theorem 3.2, this shows that $A^{t}$ is WEF-able for every round $t\in T$.
Let $A^{t}$ the allocation computed by Algorithm 1 at iteration $t$. Let $C$ be any directed cycle in $G_{A^{t},w}$, and denote $C=(i_{1},...,i_{r})$. To simplify notation, we consider $i_{1}$ as $i_{r+1}$.
We construct another allocation $B^{t}$ randomly as follows: for each agent $i_{j}\in C$, we choose one item $o_{i_{j+1}}^{t}$ uniformly from $i_{j+1}$’s bundle and transfer it to $i_{j}$’s bundle  [^1] Recall that at each iteration, each agent $i_{j}$ receives exactly $w_{i_{j}}$ items.. The expected value of $v_{i_{j}}(o_{i_{j+1}}^{t})$, the value of the item removed from $i_{j}$’s bundle, can be computed as the average value of all items in $A^{t}_{i_{j}}$: $\frac{\sum_{o\in A^{t}_{i_{j}}}v_{i_{j}}(o)}{w_{i_{j}}}=\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}$. Similarly, the expected value of $v_{i_{j}}(o_{i_{j}}^{t})$, the value of the item added the $i_{j}$’s bundle, is $\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}$. Thus, the expected change in value between $B^{t}$ and $A^{t}$ is
@@
$\sum_{i_{j}\in C}\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}$
This is exactly the total cost of cycle $C$. According to Algorithm 1, $A^{t}$ maximizes the total value among all allocations in which each agent $i$ receives exactly $w_{i}$ items. Therefore, the left-hand side of the above expression must be at most 00. But the right-hand side of the same expression is exactly the total cost of $C$. Therefore,
@@
Next, we show that the weighted envy between agents is at most $V$.
Proposition 3.11.
Let $A$ be the allocation computed by Algorithm 1. For every $i,j\in N$ it holds that $\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq V$

### v4

@@ 3.1 A given allocation: the weighted-envy graph
When trying to extend Section 2 to the weighted setup, a notable distinction arises: in the unweighted context, an allocation $A$ is EF-able if and only if $A$ maximizes the utilitarian welfare across all reassignments of its bundles to agents (part (2) of Section 2).
@@
Contrastingly, in the weighted setup, this assertion does not necessarily hold, as demonstrated by Example 1. Moreover, the converse of (3) in Section 2 can fail too. {proposition} There exists an allocation $A$, which is WEF-able, yet maximizes neither the utilitarian social welfare or the weighted utilitarian social welfare.
@@
It turns out that two out of three parts of this theorem still hold in the weighted setting; The proof is similar to the one in [halpern2019fair]. {theorem} For an allocation $A$, the following statements are equivalent:
@@
$=\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}+\frac{s_{j}}{w_{j}}.$
@@
Section 3.1 presents an effective method for verifying whether a given allocation $A$ is WEF-able. {proposition} Given an allocation $A$, it can be checked in polynomial time whether $A$ is WEF-able.
Proof.
According to Section 3.1, determining whether $A$ is is WEF-able is equivalent to verifying the absence of positive-cost cycles in the weighted envy-graph $G_{A,w}$. This can be achieved by transforming the graph by negating all edge weights and then checking for the presence of negative-cost cycles. Using the Floyd-Warshall algorithm (weisstein2008floyd, wimmer2017floyd) on the graph obtained by negating all edge cost in $G_{A,w}$ requires $O(n^{3})$ time. Constructing the initial graph $G_{A,w}$ takes $O(mn)$ time, resulting in an overall complexity of $O(mn+n^{3})$. ∎
We use Section 3.1 to present an alternative proof that part 2 of Section 2 does not hold in the weighted setting.
{proposition}
@@
As in Example 1, consider 2 bundles $B_{1},B_{2}$ and 2 agents, $N=\{i_{1},i_{2}\}$, with weights $w_{1}=1,w_{2}=10$ and valuation functions
@@
By Section 3.1, both permutations are not WEF-able. ∎
Deciding that an allocation $A$ is WEF-able is not enough. As in halpern2019fair, we also need to find a minimal subsidy vector, $\mathbf{s}$, that ensures WEF for $(A,\mathbf{s})$. This subsidy vector can be computed in strongly polynomial time and matches the one constructed in the proof of Section 3.1. {theorem} For any allocation $A$ that is WEF-able, let $\mathbf{s}^{*}$ be a subsidy vector defined by
$\mathbf{s}^{*}_{i}:=w_{i}\cdot cost_{A}(\ell_{i}(A)),$
for all $i\in N$. Then
@@
    $\Rightarrow\frac{s_{i}}{w_{i}}\geq\frac{s_{i}^{*}}{w_{i}}+\frac{s_{i_{r}}}{w_{i_{r}}}\geq\frac{s_{i}^{*}}{w_{i}}.$
@@
    The computation of $s^{*}$ can be executed as follows: Initially, the Floyd-Marshall algorithm (weisstein2008floyd, wimmer2017floyd) is applied to the graph derived by negating all edge costs in $G_{A,w}$ (This has a linear time solution since there are no cycles with positive costs in the graph). Hence, determining the longest path cost between any two agents, accomplished in $O(nm+n^{3})$ time. Subsequently, the longest path starting at each agent is identified in $O(n^{2})$ time.
@@
    Assume, for the sake of contradiction, that $s^{*}_{i}>0$ for every agent $i\in N$, which implies that $\text{cost}_{A}(\ell_{i}(A))>0$ since $w_{i}\geq 1$. Because the number of agents is finite, there must be $k\geq 2$ agents, say $i_{1},\dots,i_{k}$, such that $\ell_{i_{1}}(A),\dots,\ell_{i_{k}}(A)$ form a cycle with positive cost in $G_{w,A}$, contradicting Section 3.1.
@@
Now we can find the minimum subsidy needed in the worst-case scenario for agents with different entitlements, whether the allocation is given or can be chosen. The proofs extend those of halpern2019fair. Note that Section 3.1 applies to any additive valuations, including identical and binary cases.
{theorem}
For every weight vector and every given WEF-able allocation $A$, letting $s_{i}:=s^{*}_{i}=w_{i}\ell_{i}(A)$, the total subsidy $\sum_{i\in N}s_{i}$ is at most $\left(\frac{W}{w_{1}}-1\right)mV$, and this bound is tight in the worst case.
@@
The proofs extend those of [halpern2019fair]. By Section 3.1, to bound the subsidy required for $i$, we bound the highest cost of a path starting at $i$. We prove that, for every WEF-able allocation $A$ and agent $i$, the highest cost of a path from $i$ in $G_{A,w}$ is at most $\frac{mV}{w_{1}}$.
For every path $P$ in $G_{A,w}$,
$cost_{A}(P)=\sum_{(i,j)\in P}cost_{A}(i,j)=\sum_{(i,j)\in P}\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}$
$\leq\sum_{(i,j)\in P}\frac{v_{i}(A_{j})}{w_{1}}\leq\sum_{(i,j)\in P}\frac{V\cdot|A_{j}|}{w_{1}}\leq\frac{mV}{w_{1}}.$
Therefore, the cost of every path is at most $\frac{mV}{w_{1}}$, so agent $i$ needs a subsidy of at most $w_{i}\frac{mV}{w_{1}}$. By part (4) of Section 3.1, at least one agent has a subsidy of 0. This implies a total subsidy of at most $\frac{W-w_{1}}{w_{1}}mV=\left(\frac{W}{w_{1}}-1\right)mV$.
To establish tightness, consider an instance with $m$ identical items, which all $n$ agents value at $V$. Consider the allocation $A$, which assigns all items to a single agent $j$ with the minimum entitlement. It’s evident that $A$ is WEF-able, and its optimal subsidy vector $\mathbf{s}$ satisfies $s_{j}=0$ and $s_{i}=\frac{w_{i}}{w_{j}}mV=\frac{w_{i}}{w_{1}}mV$ for $i\neq j\in N$. Therefore, we require
$\frac{W-w_{1}}{w_{1}}mV=\left(\frac{W}{w_{1}}-1\right)mV.$
In the unweighted case $W/w_{1}=n$, so the upper bound on the subsidy becomes $(n-1)mV$. This is the same upper bound proved by halpern2019fair for the unweighted case and additive valuations.
Halpern and Shah proved in [halpern2019fair] that, if an allocation $A$ is EF-able and EF1, then every path in $G_{A,w}$ has cost at most $(n-1)V$, and therefore a subsidy of at most $(n-1)^{2}V$ is sufficient. The following theorem generalizes this result to the weighted setting. {theorem} Let $A$ be both WEF-able and $WEF(x,y)$ for any $x,y\in[0,1]$. Then there exists an envy-eliminating subsidy vector with total subsidy at most $\left(\frac{W}{w_{1}}-1\right)(n-1)V$.
@@
$\frac{v_{i}(A_{j})}{w_{j}}-\frac{v_{i}(A_{i})}{w_{i}}\leq\frac{yv_{i}(B)}{w_{i}}+\frac{xv_{i}(B)}{w_{j}}\leq$
$\frac{yv_{i}(B)}{w_{1}}+\frac{xv_{i}(B)}{w_{1}}=\frac{(x+y)v_{i}(B)}{w_{1}}\leq$
$(x+y)\frac{V}{w_{1}}.$
Any path contains at most $n-1$ arcs, that is,
$s_{i}=w_{i}\ell_{i}\leq(x+y)\frac{w_{i}}{w_{1}}(n-1)V.$
From part (4) of Section 3.1, there is at least one agent that requires no subsidy, so the required total subsidy is at most
$\frac{W-w_{1}}{w_{1}}(n-1)V=\left(\frac{W}{w_{1}}-1\right)(n-1)V.$
@@ 3.2 Computing a WEF-able allocation
We start with a lower bound on the required subsidy for general additive valuations. To prove the lower bound, we need a Lemma. {lemma} Suppose there are $n$ agents, and only one item $o$ which the agents value positively. Then an allocation is WEF-able iff $o$ is given to an agent $i$ with the highest $v_{i}(o)$.
@@
By Section 3.1, it is sufficient to check the cycles in the weighted envy-graph. If $o$ is given to $i$, then $i$’s envy is $-\frac{v_{i}(o)}{w_{i}}$, and the envy of every other agent $j$ in $i$ is $\frac{v_{j}(o)}{w_{i}}$. All other envies are 0. The only potential positive-weight cycles are cycles of length 2 involving agent $i$. The weight of such a cycle is positive iff $\frac{v_{j}(o)}{w_{i}}-\frac{v_{i}(o)}{w_{i}}>0$, which holds iff $v_{j}(o)>v_{i}(o)$. Therefore, there are no positive-weight cycles iff $v_{i}(o)$ is maximum. ∎
{theorem}
For every weight vector $\mathbf{w}$ and any $n\geq 2$, no algorithm can guarantee a total subsidy smaller than $\left(\frac{W}{w_{1}}-1\right)V$.
@@
By Section 3.2, the only WEF-able allocation is to give $o$ to agent $1$. In this case, the minimum subsidy is $s_{1}=0$ and $s_{i}=w_{i}\frac{v_{i}(o)}{w_{1}}=w_{i}V_{2nd}$. Summing all subsidies leads to $\sum_{i\geq 2}w_{i}V_{2nd}=(W-1)V_{2nd}$. As $\epsilon$ can be arbitrarily small, we get a lower bound of $(W-1)V$. ∎
In Section 1, we showed that the iterated-maximum-matching algorithm [brustle2020one] might produce an allocation that is not WEF-able.
We now introduce a new algorithm, Algorithm 1, which extends the iterated-maximum-matching approach to the weighted setting, assuming all weights are integers. The algorithm finds a one-to-many maximum matching between agents and items, ensuring that each agent $i\in N$ receives exactly $w_{i}$ items. If the number of items remaining in a round is less than $W$, we add dummy items (valued at 00 by all agents) so that the total number of items becomes $W$.
In Section 1, we add 9 dummy items, and perform a one-to-many maximum-value matching between agent and items, resulting in a WEF-able allocation: $A_{1}=\emptyset,A_{2}=\{o_{1},o_{2}\}$.
The algorithm runs in $\lceil m/W\rceil$ rounds. In each round $t$, the algorithm computes a one-to-many maximum-value matching $\{A_{i}^{t}\}_{i\in N}$ between all agents and unallocated items $O_{t}$, where each agent $i\in N$ receives exactly $w_{i}$ items.
To achieve this, we reduce the problem to the minimum-cost network flow problem (goldberg1989network) by constructing a flow network and computing the maximum integral flow of minimum cost. The flow network is defined as follows:
  • Layer 1 (Source Node). a single source node $s$.
  • Layer 2 (Agents). a node for each agent $i\in N$, with an arc from $s$ to $i$, having cost 00 and capacity $w_{i}$.
  • Layer 3 (Unallocated Items). a node for each unallocated item $o\in O_{t}$, with an arc from each agent $i\in N$ to item $o$, having cost $-v_{i}(o)$ and capacity $1$.
  • Layer 4 (Sink Node). a single sink node $t$, with an arc from each item $o\in O_{t}$ to $t$, having 0 cost and capacity $1$.
Any integral maximum flow in this network corresponds to a valid matching where each agent $i\in N$ receives exactly $w_{i}$ items from $O_{t}$, and each item is assigned to exactly one agent, the result is a minimum-cost one-to-many matching based on the costs in the constructed network. Because we negate the original costs in our construction, the obtained matching $\{A_{i}^{t}\}_{i\in N}$ maximizes the total value with respect to the original valuations. After at most $\lceil m/W\rceil$ valuations, all items are allocated.
@@
Algorithm 1 Weighted Sequence Protocol For Additive Valuations and Integer weights
1:Instance $(N,M,v,\mathbf{w})$ with additive valuations.
2:WEF-able allocation $A$ with total required subsidy of at most $(W-w_{1})V$.
@@
5:while $O_{t}\neq\emptyset$ do Construct the flow network $G^{\prime}=(V^{\prime},E^{\prime})$:
  • define $V^{\prime}=N\cup O_{t}\cup\{s,t\}$.
  • Add arcs with the following properties:
    • From $s$ to each agent $i\in N$ with cost 00 and capacity $w_{i}$.
    • From each agent $i\in N$ to each unallocated item $o\in O_{t}$, with cost $-v_{i}(o)$ and capacity 1.
    • From each unallocated item $o\in O_{t}$ to $t$ with cost 00 and capacity $1$.
Compute an integral maximum flow of minimum cost on $G^{\prime}$, resulting in the one-to-many matching $\{A_{i}^{t}\}_{i\in N}$
@@
7:     $t\leftarrow t+1$
8:end while
9:return $A$
{proposition}
@@
We prove that, in every round $t$, the total cost added to any directed cycle in the weighted-envy graph is non-positive. Combined with Section 3.1, this shows that $A^{t}$ is WEF-able for every round $t\in T$.
Let $A^{t}$ the allocation computed by Algorithm 1 at iteration $t$. Note that Algorithm 1 is deterministic. Let $C$ be any directed cycle in $G_{A^{t},w}$, and denote $C=(i_{1},...,i_{r})$. To simplify notation, we consider $i_{1}$ as $i_{r+1}$.
Given the allocation $A^{t}$ and the cycle $C$, we construct a random alternative allocation $B^{t}$ as follows: for each agent $i_{j}\in C$, we choose one item $o_{i_{j+1}}^{t}$ uniformly from $i_{j+1}$’s bundle and transfer it to $i_{j}$’s bundle  [^2] Recall that at each iteration, each agent $i_{j}$ receives exactly $w_{i_{j}}$ items.. The expected value of $v_{i_{j}}(o_{i_{j+1}}^{t})$, the value of the item removed from $i_{j}$’s bundle, can be computed as the average value of all items in $A^{t}_{i_{j}}$:
$\frac{\sum_{o\in A^{t}_{i_{j}}}v_{i_{j}}(o)}{w_{i_{j}}}=\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}.$
Similarly, the expected value of $v_{i_{j}}(o_{i_{j}}^{t})$, the value of the item added the $i_{j}$’s bundle, is $\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}$. Thus, the expected change in value between $B^{t}$ and $A^{t}$ is
@@
$\sum_{i_{j}\in C}\frac{v_{i_{j}}(A^{t}_{i_{j+1}})}{w_{i_{j+1}}}-\frac{v_{i_{j}}(A^{t}_{i_{j}})}{w_{i_{j}}}.$
This is exactly the total cost of cycle $C$. According to Algorithm 1, $A^{t}$ maximizes the total value among all allocations in which each agent $i$ receives exactly $w_{i}$ items. Therefore, the left-hand side of the above expression, which is the difference between the sum of values in $B^{t}$ and the sum of values in $A^{t}$, must be at most 00. But the right-hand side of the same expression is exactly the total cost of $C$. Therefore,
@@
As the allocation in each iteration is WEF-able, the output allocation $A$ is WEF-able too. To compute an upper bound on the subsidy, we adapt the proof technique in [brustle2020one]. Let $A^{t}$ be the output allocation from Algorithm 1, computed in iteration $t$. For each $i\in N$, we define the modified valuation function as follow:
$\bar{v}_{i}(A_{j}^{t})=\begin{cases}\frac{v_{i}(A_{i}^{t})}{w_{i}}&j=i\\ \frac{v_{i}(A_{j}^{T})}{w_{j}}&j\neq i,t=T\\ \max{\left(\frac{v_{i}(A_{j}^{t})}{w_{j}},\frac{v_{i}(A_{i}^{t+1})}{w_{i}}\right)}&j\neq i,t<T\end{cases}$
Under the modified valuations, for any two agents $i,j\in N$, the modified-cost assigned to the edge $(i,j)$ in the envy graph (with unit weights) is defined as $\overline{cost}_{A}(i,j)=\bar{v}_{i}(A_{i})-\bar{v}_{i}(A_{j})$. Moreover, the modified-cost of a path $(i_{1},...,i_{k})$ is $\overline{cost}_{A}(i_{1},...,i_{k})=\sum_{j=1}^{k-1}\overline{cost}_{A}(i_{j},i_{j+1})$.
{lemma}
Let $A$ be a WEF-able allocation. For any positive number $z$, if $cost_{A}(i,k)\geq-z$ for every edge $(i,k)$ in $G_{A,w}$, then the maximum subsidy required is at most $w_{i}z$ per agent $i\in N$.